# Fading in the Flow: Suppression of cold gas growth in expanding galactic outflows

### v2

@@ 4.1.1 Homogeneous cloud-crushing simulations
Figure 9: Evolution of the lateral extent of a cloud (in units of the initial cloud radius) as a function of the distance travelled by the center of mass of the cloud (normalized by the initial cloud distance from the center of the CC85 wind). The cloud increases in its lateral extent as it moves downstream. Along each of the curves, the diamonds denote the instant when the clouds attain twice their initial mass and the stars mark the instant when the center of mass of the clouds are travelling at $80$ percent of the local wind speed – the clouds have enterned an ‘entrained’ stage. Once the clouds are entrained, we expect the geometry of the CC85 wind to make the clouds spread proportional to the distance they travel, as indicated by the yellow band. Most of the clouds in their entrained state seem to follow this trend ($\propto d_{\rm cl}$), except for the largest clouds (small $t_{\rm cool,mx}/t_{\rm cc}|_{\rm ini}$) which are not entrained in a significant part of their journey as well as their lateral expansion is smaller compared to their initial size.
@@
Figure 10: Evolution of the cloud mass (in units of the initial cloud mass) as a function of the distance travelled by the center of mass of the cloud (normalized by the initial cloud distance from the center of the CC85 wind). As discussed in Section 4.1.1, the increase in the lateral spread of the cloud (cf. Fig. 9) results in a steeper increase in its mass $m_{\rm cl}$ with distance $d_{\rm cl}$ ($m_{\rm cl}\propto d_{\rm cl}^{23/12}\approx d_{\rm cl}^{2}$; gray band) compared to the previous estimate by Gronke & Oh 2020 ($m_{\rm cl}\propto d_{\rm cl}^{7/6}\approx d_{\rm cl}$; yellow band; GO20). Along each of the curves, the diamonds and stars have the same meaning as in Fig. 9. The clouds which have travelled a long distance after entrainment approximately follows $m_{\rm cl}\propto d_{\rm cl}^{2}$, in line with our analytic estimates (gray band). For the largest clouds, the lateral expansion is lower compared to their initial size (closer to the $t_{\rm sc,cl}\sim\rm const$ assumption of GO20; yellow band). This also leads to a different scaling relation between the cloud surface area and distance travelled by the cloud compared to GO20 (cf. Fig. 16).
@@
While a homogeneous and non-varying background is not particularly realistic in astrophysical settings, the prior studies employing this setup focused on several different physical mechanisms at play during a cloud-wind interaction. One notable difference from simulations of cloud-crushing in a homogeneous background is that, due to the expansion of the domain, the thickness of the tail increases as the cloud moves downstream (the distance of its center of mass $d_{\rm cl}$ increases). In a CC85 wind, we note that once the cloud is co-moving with the background wind, the orthogonal spread of the cloud $R_{\rm cl,\perp}$ increases with distance $\propto d_{\rm cl}$. This is illustrated in Fig. 9 which shows the lateral extent of the clouds, calculated as the difference between the $95$ and $5$ percentiles of the perpendicular distance (with respect to the wind direction) of all the cold gas ($T<8\times 10^{4}\ \rm K$; cf. Section 2.3), as a function of the position of the center of mass of the cloud. In contrast, $R_{\rm cl,\perp}$ is a constant in a uniform wind. The relation between the transverse cloud spread and the distance it travels introduces significant differences in the rate at which mass grows in entrained clouds compared to clouds in a uniform wind. We discuss using scaling relations from previous simulations of radiative turbulent mixing layers, how the orthogonal expansion of clouds can explain the supression of cold mass growth in a CC85 wind compared to ‘vanilla’ cloud-crushing (cf. Section 3.2 & Fig. 5).
@@
In the entrained state (i.e., after substantial mass and momentum exchange between the cold cloud and the background wind), the rate of mass growth of a cloud $\dot{m}_{\rm cl}\sim A_{\rm cl}\rho_{\rm wind}v_{\rm mix}$ depends on the inflow velocity into the radiative mixing layers $v_{\rm mix}$, the surface area of the cloud $A_{\rm cl}$, and the density of the wind $\rho_{\rm wind}$. We first discuss the analytic scaling relations of cloud properties and turbulent boundary layers presented in Gronke & Oh 2020 (GO20 hereafter) in light of our simulation results and then compare our findings with the simulations carried out by GO20 approximating an expanding outflow. GO20 found that the inflow velocity in the mixing layer follows $v_{\rm mix}\sim c_{\rm s,cl}(t_{\rm sc,cl}/t_{\rm cool,cl})^{1/4}$, roughly consistent with later mixing layer simulations (Fielding et al. 2020; Tan et al. 2021). GO20 assumed that the sound crossing time across the shortest dimension of the cloud remains constant. However, in a CC85 wind, we expect the cloud lateral extent to increase with distance ($R_{\rm cl,\perp}\propto d_{\rm cl}$; cf. Fig. 9), implying $t_{\rm sc,cl}\sim R_{\rm cl,\perp}/c_{\rm s,cl}\sim d_{\rm cl}/c_{\rm s,cl}$. Once the cloud is entrained, similar to GO20, we find that its surface area $A_{\rm cl}\propto V_{\rm cl}^{2/3}\sim(m_{\rm cl}/\rho_{\rm cl})^{2/3}\sim m_{\rm cl}^{2/3}d_{\rm cl}^{4\gamma/3}$ ($V_{\rm cl}$ is the cloud volume; see upper panel of Fig. 16 in Appendix B) due to the pressure drop $p_{\rm cl}\propto d_{\rm cl}^{-2\gamma}$ in a CC85 wind.
@@
Hence, the mass of a cold cloud in the entrained state $m_{\rm cl}\propto d_{\rm cl}^{5\gamma/2-9/4}\sim d_{\rm cl}^{23/12}\approx d_{\rm cl}^{2}$ (here $\gamma=5/3$). This is a significantly steeper relation with cloud distance than what is expected from GO20 that uses a constant sound crossing time – $m_{\rm cl}\propto d_{\rm cl}^{5\gamma/2-3}\sim d_{\rm cl}^{7/6}\approx d_{\rm cl}$. In Fig. 10, we demonstrate that with the inclusion of cloud expansion in the sound crossing time, we are able to explain the steeper scaling (compared to the analytic arguments in GO20) of cold mass with distance travelled by clouds that are entrained for a long time. [^5] This inclusion of cloud expansion, however, weakly alters the scaling of cloud surface area with the distance travelled as the cloud traverses downstream (see Appendix B; lower panel of Fig. 16).
@@
Figure 11: Comparison of the cold mass evolution of gas from our cloud-crushing simulation in CC85 wind with the cloud-crushing simulation employing an approximate expanding box model from Gronke & Oh 2020 (GO20; see Section 4.1.1 & Appendix C for details). For this comparison, we re-implemented the expanding box model used by GO20 in PLUTO and re-simulated the cloud-crushing problem using our fiducial parameters, i.e., $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$ and an initial Mach number $\mathcal{M}=1.496$ in the wind (cf. Table 1). In both the simulations, the cloud radius is initially resolved by $16$ grid cells. Follow this link to see the evolution of this cloud in the expanding box simulation.
@@
Moving beyond analytic scaling estimates, we now directly compare simulation results from GO20 with ours. Specifically, GO20 studied the evolution of cold clouds in expanding outflows by employing a model of an expanding box in their simulations that approximates the azimuthal and polar expansion that the clouds undergo as they move downstream due to the geometry of the wind (also, see, similar setups by Heitsch & Putman, 2009; Tan et al., 2023 for infalling clouds). This model employs an anisotropic expansion in the simulation domain to model the CC85 wind which expands in the polar and azimuthal directions but not in the radial direction (cf. Dutta & Sharma 2019; also Scannapieco 2017 who erroneously assumed the wind to be isotropically expanding). This model assumes that the size of the cloud (and the simulation domain containing and following it) is always small enough compared to the length scale on which the wind property changes, and the wind can be assumed to be plane parallel. Therefore, under these assumptions, the entire cloud always experiences a uniform expansion perpendicular to the wind due to its movement downstream. Consequently, a local simulation domain that follows the cloud gets expanded in all directions perpendicular to the direction of the wind, determined by an expansion factor that depends on the instantaneous position of the cloud with respect to its initial position.
@@
In contrast, our cloud-crushing simulations in an expanding CC85 wind do not make such simplifying assumptions and can account for the differential expansion of a cloud along its tail. This differential expansion of a cloud along the wind and the curvature of the wind (i.e., variation in wind properties along the cloud) become increasingly important for large clouds deeper in the cooling regime. We find that this differential expansion can lead to a lower expansion in the cloud head compared to its tail (cf. Fig. 18 in Appendix C), which, in turn, can result in an overall increase in the tail length of the cometary cloud structures (after several $t_{\rm cc}$) compared to the uniform expansion considered by GO20. Moreover, the background wind properties – such as density, temperature, and pressure – vary significantly with the distance $d_{\rm cl}$ travelled by the cloud. Consequently, even modest discrepancies in estimating how this distance changes with time (arising from differences in how mixing/cooling between the cold cloud and hot wind progresses, which affects mass and momentum exchange) can lead to substantial differences in the predicted cold mass evolution, especially at late times.
@@
Another difference from GO20 is that we initialize our clouds at a distance where the wind Mach number is between 1-2, not too far from the sonic point. Therefore the wind is not in the asymptotic scaling ($\rho\propto d^{-2}$, $p\propto d^{-2\gamma}$) regime, an assumption made in GO20. All of these factors potentially contribute to differences in the cold mass growth between the two simulation setups, which increase as the simulations progress further in time. This is illustrated in Fig. 11, where we compare the evolution of the cold mass in the two simulation setups for our fiducial parameters ($t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$, and $\mathcal{M}=1.496$; cf. Table 1; see Appendix C for our implementation of GO20 setup). Since the cold mass estimate is slightly lower in the anisotropic expanding box of GO20 compared to our simulation domain, it might explain why GO20 reported a flatter mass-distance relation ($m_{\rm cl}\propto d_{\rm cl}$). We expect the differences in the cold mass estimate between these two types of simulation setups to increase further in time.
@@ 4.1.2 Semi-analytic multiphase wind models
Fielding & Bryan 2022 (FB22 hereafter) introduced a model that describes the evolution of a cold cloud that starts with some initial velocity (or from rest) in an outflow (Section 3 in their paper). FB22 assumed a steady injection of many of such clouds, possibly lifted from the ISM, at a constant rate within a region close to the center of the wind. FB22 introduced a model for the evolution of an individual cloud in a wind and used it as a source of energy, mass, and momentum injection/removal in a steady wind, assuming that clouds are introduced at a fixed rate within a central region in an outflow. Consequently, a model to account for the evolution of a collection of small clouds continuously injected into an outflow has been utilized in developing a general framework to predict the structure of a supernovae-driven multiphase galactic outflow. Although in FB22 the single-cloud model was utilized to take into account the evolution of a collection of clouds in a spatially changing background wind (that act as source terms in the wind equation), the detailed evolution of an individual cloud in a non-uniform wind, and its comparison with simulations remain largely unexplored. [^6] Albeit see the similar work by Nikolis & Gronke 2024 that carried out a direct comparison to hydrodynamical simulations. The evolution of a cloud seeded in a uniform background from the single-cloud model has been illustrated in FB22. Since our work focuses on simulations of how individual clouds evolve in both a uniform and an expanding CC85 (Chevalier & Clegg 1985) wind, we are able to compare the FB22 model prediction of cloud-wind interation with our simulation results. Furthermore, the constraints obtained by the validation of the single-cloud model would enable the construction of realistic wind profiles by providing accurate source terms in the wind equation. Prior to a comparison with FB22, it is worth noting that our work focuses on the detailed behavior of a cloud, considering a neighborhood close to the cloud. The back reaction of the cloud on the volume-filling gas that comprises the wind is beyond the scope of this work. In addition, we are mostly interested in modeling the evolution of an individual cloud and do not consider the collective impact of multiple clouds that can potentially alter the large-scale properties of the wind.
@@
Figure 12: Comparison of the mass evolution of gas from our cloud-crushing simulation with that predicted from the FB22 single-cloud model for ‘vanilla’ cloud-crushing (in uniform wind) and for cloud-crushing in a CC85 wind (see Section 4.1.2 for details). The dashed lines are the simulation results for ‘vanilla’ cloud-crushing (in orange) and cloud-crushing in a CC85 background (in blue). As a specific example which would clearly demonstrate departure from model prediction, we choose the simulation result with $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.8$ and $\mathcal{M}=1.496$ (cf. Table 1; not the fiducial parameter set). The corresponding model predictions are shown with solid lines in the same color as their dashed counterparts. The solid light colored lines use the original prescription from FB22 and does not match well with the simulation results. With the tuned parameters (described in Section 4.1.2 & Appendix E), we are able to achieve a better match with the simulation results – the solid dark colored lines. For simplicity, the free parameters of the model are chosen by visual inspection and are listed as follows, ($f_{\rm cool},f_{\rm mix},f_{\rm turb},C_{\rm drag},\xi_{\rm trans}$) = ($2.0,2.1,0.08,0.5,13$). Additionally, $f_{\rm dim}(\xi<\xi_{\rm trans})=0.49$ and $f_{\rm dim}(\xi\geq\xi_{\rm trans})$ equals $0.70$ for vanilla cloud-crushing and $0.68$ for cloud-crushing in CC85 wind. In this context, we simply demonstrate the need to construct a robust single-cloud model – a necessary step to explain the evolution of clouds in realistic outflows (cf. Fig. 13).
To facilitate a comparison with the cloud evolution model and the results from our simulations, we consider the single-cloud model introduced by FB22 (Section 3 in their paper). [^7] However, for simplicity and also to compare with our simulations, we do not consider different metallicities for the cloud and the wind. On comparing the original FB22 cloud-wind model with our simulation results, we find an unsatisfactory model prediction for the evolution of cold mass growth in a CC85 wind (see the discussion on Fig. 12 later in this section for details). Therefore, we modify the model presented by FB22 to better reproduce the evolution of the cold mass from our simulations. Next, we briefly discuss the single-cloud model by FB22. It is an empirical model of the evolution of a cloud in embedded in a hot wind. We also discuss the modifications we introduce to the parameters of the original model in order to obtain a significantly improved match with our simulation results. In the remainder of this section (and Appendix E), we describe the single-cloud model, our modifications to it, and also demonstrate how well this model compares against our simulation.
The single-cloud model introduced by FB22 is a set of coupled ordinary differential equations which can be numerically solved with some initial condition (described elaborately in Appendix E). The most important part in this model is incorporating radiative cooling of the turbulent mixing layers as the cloud moves and mixes in the wind. This essentially captures the growth/destruction of cold clouds in the wind, and it is taken into account by the factors $\xi$ and $\alpha$ in Eq. E.1. The key term in the model determining the cold mass evolution in the wind, is the factor $\alpha$ that appears as an exponent on $\xi$. In physical terms, $\alpha$ depends on the fractal dimension of the radiatively cooling turbulent mixing layer formed between the cloud and the wind due to shear-driven hydrodynamic instabilities (Mandelbrot 1975; Federrath et al. 2009; Fielding et al. 2020; Tan & Fielding 2024). Since the factor $\alpha$ appears in the exponent of $\xi$, the rate at which cold gas forms in the FB22 model can change by orders of magnitude due to even minute changes in $\alpha$.
In addition to this, simulations (Ji et al. 2019; Tan et al. 2021) show that there is a transition in the mass growth rate due to radiative cooling from the slow cooling regime (small values of $\xi$) to the fast cooling regime (large values of $\xi$) at $\xi_{\rm trans}$ (a free parameter in the model). According to the model introduced by Fielding et al. 2020; Fielding & Bryan 2022, this transition corresponds to a change in the fractal dimension factor $f_{\rm dim}$ causing a change in the value of $\alpha=(3f_{\rm dim}-1)/2$. Here $f_{\rm dim}$ follows the usual definition of fractal dimension, such that the area of a fractal surface $A_{\Delta}$ depends on the measurement scale $\Delta$, and $A_{\Delta}/L^{2}=(L/\Delta)^{f_{\rm dim}}$, where $L$ is the linear size of the fractal surface (Sreenivasan et al. 1989). Specifically, FB22 use $\alpha=0.25$ for the strong cooling ($\xi>1$) and $\alpha=0.5$ for the weak cooling ($\xi<1$) regime (see Section 4.3 in Tan et al. 2021 for a detailed discussion on the choice of these values).
Fig. 12 shows a comparison of our numerical simulations – both in an expanding CC85 (blue lines) as well as uniform plane-parallel wind (orange lines) – with the prediction of single-cloud evolution from the FB22 model (shown in solid light color in Fig. 12). One can see that the numerical results do not match the semi-empirical model sufficiently, with $\gtrsim 1$ dex more cold gas mass in the simulations compared to the model. The FB22 model is also unable to reproduce the suppression of cold mass growth of clouds in an expanding wind when compared to a uniform plane-parallel wind – a definitive trend in our simulation. This is concerning, since the FB22 is used as a basis for multiphase subgrid model of galactic winds (Smith et al. 2024a, b) with the aim of being eventually incorporated into cosmological simulations. This mistmatch in the cold gas mass also propagates to the overall wind properties – as we discuss below.
We also show in Fig. 12, using solid dark colored curves, a new model prediction with tuned FB22 parameters in order to reproduce the simulation data. Specifically, we changed the transition from strong to weak cooling from $\xi_{\rm trans}=1$ to $\xi_{\rm trans}=13$, and more importantly, we adjusted $\alpha$ to $0.235$ for $\xi<\xi_{\rm trans}$ and to $0.55$ [$0.52$] for $\xi>\xi_{\rm trans}$ in the uniform [expanding] wind. Crucially, this implies that mass growth, which depends very sensitively on $\alpha$, is drastically different across different wind properties. Note, that while we managed to reproduce the simulation runs shown in Fig. 12 using these values, [^8] Note that although the $\alpha$ values that can reproduce the evolution shown in Fig. 12 corresponds to a fractal dimension of $0.49$ in the weak cooling regime and $0.7[0.68]$ in the strong cooling regime in a uniform [expanding] wind. These values lack sound physical justification, pointing potentially to inadequacies of the model., they are not robust across wind or cloud properties and would need to be adjusted depending on the wind+cloud parameters. We find that such a modification is necessary to explain the trend in the evolution of the cold mass from our simulations, especially to explain the suppression of cold mass growth in a CC85 wind compared to a uniform one. It should be noted that the FB22 model considers all gas at exactly the cloud temperature $T_{\rm cl}$ to be cold, whereas in our simulations, all gas cells with temperature lower than $2T_{\rm cl}$ are considered cold. This allows us to avoid testing equality of floating-point numbers. Nevertheless, because the cloud temperature has a very tight spread about $T_{\rm cl}$ (cf. Section 3.5), our cold phase identification scheme to compare with FB22 model prediction is relatively robust. It is important to emphasize that we do not claim the parameters or modifications to the FB22 model presented here to be a best-fit solution. A detailed exploration of the model’s parameter space is beyond the scope of this work and is left for future studies. Our primary goal is to emphasize the need for improving existing models by calibrating them against cloud-crushing simulations in evolving background wind conditions.

### v3

@@ 4.1.1 Homogeneous cloud-crushing simulations
Figure 10: Evolution of the lateral extent of a cloud $R_{\rm cl,\perp}$ (in units of the initial cloud radius) as a function of the distance travelled by the centre of mass of the cloud (normalised by the initial cloud distance from the centre of the CC85 wind). The lateral extent of the cloud has been calculated as the difference between the $95$ and 55 percentile values of the perpendicular distance of the cloud material from the axis passing through the centre of the cloud, parallel to the wind. The cloud increases in its lateral extent as it moves downstream. Along each of the curves, the diamonds denote the instant when the clouds attain twice their initial mass and the stars mark the instant when the centre of mass of the clouds are travelling at $80$ percent of the local wind speed – the clouds have entered an ‘entrained’ stage. Once the clouds are entrained, we expect the geometry of the CC85 wind to make the clouds spread proportional to the distance they travel, as indicated by the yellow band. Most of the clouds in their entrained state seem to follow this trend ($\propto d_{\rm cl}$), except for the largest clouds (small $t_{\rm cool,mx}/t_{\rm cc}|_{\rm ini}$) which are not entrained in a significant part of their journey as well as their lateral expansion is smaller compared to their initial size. Lateral expansion in these largest clouds is, therefore, not determined by the (CC85) wind geometry but by their isobaric expansion with the wind following a much steeper relation $R_{\rm cl,\perp}\propto d_{\rm cl}^{\gamma}$, indicated by the cyan band and discussed in detail in the text.
@@
Figure 11: Evolution of the cloud mass (in units of the initial cloud mass) as a function of the distance travelled by the centre of mass of the cloud (normalised by the initial cloud distance from the centre of the CC85 wind). As discussed in Section 4.1.1, the increase in the lateral spread of the cloud (cf. Fig. 10) results in a steeper increase in its mass $m_{\rm cl}$ with distance $d_{\rm cl}$ ($m_{\rm cl}\propto d_{\rm cl}^{23/12}\approx d_{\rm cl}^{2}$; cyan band) compared to the previous estimate by Gronke & Oh 2020 in a CC85 wind ($m_{\rm cl}\propto d_{\rm cl}^{7/6}\approx d_{\rm cl}$; yellow band; GO20). Along each of the curves, the stars have the same meaning as in Fig. 10. The clouds which have travelled a long distance after entrainment approximately follows $m_{\rm cl}\propto d_{\rm cl}^{2}$, in line with our analytic estimate (cyan band). For the largest clouds, the lateral expansion is lower compared to their initial size (closer to the $t_{\rm sc,cl}\sim\rm const$ assumption of GO20; yellow band). This also leads to a different scaling relation between the cloud surface area and distance travelled by the cloud in a CC85 wind compared to GO20 (cf. Fig. 17). Note that both of these mass scaling relations with distance are flatter than what is expected in a uniform wind ($m_{\rm cl}\propto d_{\rm cl}^{3}$) – explaining the reduced cold mass formation in an expanding wind.
@@
While a homogeneous and non-varying background is not particularly realistic in astrophysical settings, the prior studies employing this setup focused on several different physical mechanisms at play during a cloud-wind interaction. One notable difference from simulations of cloud-crushing in a homogeneous background is that, due to the expansion of the domain, the thickness of the tail increases as the cloud moves downstream (the distance of its centre of mass $d_{\rm cl}$ increases). In a CC85 wind, we note that once the cloud is co-moving with the background wind, the orthogonal spread of the cloud $R_{\rm cl,\perp}$ increases with distance $\propto d_{\rm cl}$. This is illustrated in Fig. 10 which shows the lateral extent of the clouds, calculated as the difference between the $95$ and $5$ percentiles of the perpendicular distance (with respect to the wind direction) of all the cold gas ($T<8\times 10^{4}\ \rm K$; cf. Section 2.3), as a function of the position of the centre of mass of the cloud. In contrast, $R_{\rm cl,\perp}$ is a constant in a uniform wind. The relation between the transverse cloud spread and the distance it travels introduces significant differences in the rate at which mass grows in entrained clouds compared to clouds in a uniform wind. We discuss using scaling relations from previous simulations of radiative turbulent mixing layers, how the orthogonal expansion of clouds can explain the suppression of cold mass growth in a CC85 wind compared to ‘vanilla’ cloud-crushing (cf. Section 3.2 & Fig. 5).
@@
In the entrained state (i.e., after substantial mass and momentum exchange between the cold cloud and the background wind), the rate of mass growth of a cloud $\dot{m}_{\rm cl}\sim A_{\rm cl}\rho_{\rm wind}v_{\rm mix}$ depends on the inflow velocity into the radiative mixing layers $v_{\rm mix}$, the surface area of the cloud $A_{\rm cl}$, and the density of the wind $\rho_{\rm wind}$. Once the cloud is entrained, its velocity is almost constant and close to the asymptotic wind velocity $v_{\rm wind,\infty}$. We first discuss the analytic scaling relations of cloud properties and turbulent boundary layers presented in Gronke & Oh 2020 (GO20 hereafter) in light of our simulation results, and then compare our findings with the simulations carried out by GO20 approximating an expanding outflow. GO20 found that the inflow velocity in the mixing layer follows $v_{\rm mix}\sim c_{\rm s,cl}(t_{\rm sc,cl}/t_{\rm cool,cl})^{1/4}$, roughly consistent with later mixing layer simulations (Fielding et al. 2020; Tan et al. 2021). GO20 assumed that the sound crossing time across the shortest dimension of the cloud remains constant. However, in a CC85 wind, we expect the clouds, once entrained, to follow the wind geometry and have their lateral extent to increase with distance ($R_{\rm cl,\perp}\propto d_{\rm cl}$; cf. Fig. 10), [^5] This scaling of the lateral extent of a cloud, proportional to the distance travelled by it down the wind, is also the basis of the anisotropic expanding box setup used by GO20 and discussed in Section 4.1.1 & Appendix D. implying $t_{\rm sc,cl}\sim R_{\rm cl,\perp}/c_{\rm s,cl}\sim d_{\rm cl}/c_{\rm s,cl}$. Once the cloud is entrained, similar to GO20, we find that its surface area $A_{\rm cl}\propto V_{\rm cl}^{2/3}\sim(m_{\rm cl}/\rho_{\rm cl})^{2/3}\sim m_{\rm cl}^{2/3}d_{\rm cl}^{4\gamma/3}$ ($V_{\rm cl}$ is the cloud volume; see upper panel of Fig. 17 in Appendix B) due to the pressure drop $p_{\rm cl}\propto d_{\rm cl}^{-2\gamma}$ in a CC85 wind.
@@
Hence, the mass of a cold cloud in the entrained state $m_{\rm cl}\propto d_{\rm cl}^{5\gamma/2-9/4}\sim d_{\rm cl}^{23/12}\approx d_{\rm cl}^{2};\gamma=5/3$. This is a significantly steeper relation with cloud distance than what is expected from GO20 that uses a constant sound crossing time – $m_{\rm cl}\propto d_{\rm cl}^{5\gamma/2-3}\sim d_{\rm cl}^{7/6}\approx d_{\rm cl};\gamma=5/3$). Consequently, post entrainment ($A_{\rm cl}\propto V_{\rm cl}^{2/3}$), because of geometric the expansion orthogonal to the wind, small clouds follow $m_{\rm cl}\propto d_{\rm cl}^{5\gamma/2-9/4}$, which results in $A_{\rm cl}\propto d_{\rm cl}^{3(\gamma-1/2)}$. In contrast, for the largest clouds, undergoing lesser orthogonal expansion compared to their initial size, the sound crossing time across them remain constant (such that $m_{\rm cl}\propto d_{\rm cl}^{5\gamma/2-3}$), and their surface area $A_{\rm cl}\propto d_{\rm cl}^{3\gamma-2}$ (see lower panel of Fig. 17 in Appendix B). In Fig. 11, we demonstrate that with the inclusion of cloud expansion in the sound crossing time, we are able to explain the steeper scaling (compared to the analytic arguments in GO20) of cold mass with distance travelled by clouds that are entrained for a long time. [^6] This inclusion of cloud expansion, however, weakly alters the scaling of cloud surface area with the distance travelled as the cloud traverses downstream (see Appendix B; lower panel of Fig. 17). We can also explain the steeper lateral extent with distance travelled for the largest clouds (cf. Fig. 10) by noting that for these large clouds the expansion of the CC85 background compared to their size is of lesser importance. Instead, as clouds maintain pressure equilibrium with the background, $R_{\rm cl,\perp}\propto\rho_{\rm cl}^{-1/2}(m_{\rm cl}/R_{\rm cl,\parallel})^{1/2}\sim d_{\rm cl}^{\gamma}$ – a steeper scaling with distance for the large clouds (cyan band in Fig. 10). This scaling follows from the fact that for the largest clouds $m_{\rm cl}\propto d_{\rm cl}$ (cf. Fig. 11) and they are not sufficiently entrained in the wind to follow the wind expansion. In addition, due to being not entrained, their length along the wind $R_{\rm cl,\parallel}\propto d_{\rm cl}$ (as expected from ‘pure stretching’) – keeps growing and does not saturate to a fixed value unlike the smaller clouds (see Fig. 18 for the $R_{\rm cl,\parallel}(d_{\rm cl})$ evolution).
@@
Figure 12: Comparison of the cold mass evolution of gas from our cloud-crushing simulation in CC85 wind with the cloud-crushing simulation employing an approximate expanding box model from Gronke & Oh 2020 (GO20; see Section 4.1.1 & Appendix D for details). For this comparison, we re-implemented the expanding box model used by GO20 in PLUTO and re-simulated the cloud-crushing problem using our fiducial parameters, i.e., $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$ and an initial Mach number $\mathcal{M}=1.496$ in the wind (cf. Table 1). In both the simulations, the cloud radius is initially resolved by $16$ grid cells. Follow this link to see the evolution of this cloud in the expanding box simulation.
@@
Moving beyond analytic scaling estimates, we now directly compare simulation results from GO20 with ours. Specifically, GO20 studied the evolution of cold clouds in expanding outflows by employing a model of an expanding box in their simulations that approximates the azimuthal and polar expansion that the clouds undergo as they move downstream due to the geometry of the wind (also, see, similar setups by Heitsch & Putman, 2009; Tan et al., 2023 for in-falling clouds). This model employs an anisotropic expansion in the simulation domain to model the CC85 wind which expands in the polar and azimuthal directions but not in the radial direction (cf. Dutta & Sharma 2019; also Scannapieco 2017 who erroneously assumed the wind to be isotropically expanding). This model assumes that the size of the cloud (and the simulation domain containing and following it) is always small enough compared to the length scale on which the wind property changes, and the wind can be assumed to be plane parallel. Therefore, under these assumptions, the entire cloud always experiences a uniform expansion perpendicular to the wind due to its movement downstream. Consequently, a local simulation domain that follows the cloud gets expanded in all directions perpendicular to the direction of the wind, determined by an expansion factor that depends on the instantaneous position of the cloud with respect to its initial position.
@@
In contrast, our cloud-crushing simulations in an expanding CC85 wind do not make such simplifying assumptions and can account for the differential expansion of a cloud along its tail. This differential expansion of a cloud along the wind and the curvature of the wind (i.e., variation in wind properties along the cloud) become increasingly important for large clouds deeper in the cooling regime. We find that this differential expansion can lead to a lower expansion in the cloud head compared to its tail (cf. Fig. 21 in Appendix D), which, in turn, can result in an overall increase in the tail length of the cometary cloud structures (after several $t_{\rm cc}$) compared to the uniform expansion considered by GO20. Moreover, the background wind properties – such as density, temperature, and pressure – vary significantly with the distance $d_{\rm cl}$ travelled by the cloud. Consequently, even modest discrepancies in estimating how this distance changes with time (arising from differences in how mixing/cooling between the cold cloud and hot wind progresses, which affects mass and momentum exchange) can lead to substantial differences in the predicted cold mass evolution, especially at late times.
@@
Another difference from GO20 is that we initialise our clouds at a distance where the wind Mach number is between 1-2, not too far from the sonic point. Therefore, the wind is not in the asymptotic scaling ($\rho\propto d^{-2}$, $p\propto d^{-2\gamma}$) regime, an assumption made in GO20. All of these factors potentially contribute to differences in the cold mass growth between the two simulation setups, which increase as the simulations progress further in time. This is illustrated in Fig. 12, where we compare the evolution of the cold mass in the two simulation setups for our fiducial parameters ($t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$, and $\mathcal{M}=1.496$; cf. Table 1; see Appendix D for our implementation of GO20 setup). Since the cold mass estimate is slightly lower in the anisotropic expanding box of GO20 compared to our simulation domain, it might explain why GO20 reported a flatter mass-distance relation ($m_{\rm cl}\propto d_{\rm cl}$). We expect the differences in the cold mass estimate between these two types of simulation setups to increase further in time.
@@ 4.1.2 Semi-analytic multiphase wind models
Figure 13: Comparison of the cloud mass evolution from our cloud-crushing simulation with that predicted from the FB22 single-cloud model for ‘vanilla’ cloud-crushing (in uniform wind) and for cloud-crushing in an expanding CC85 wind (see Section 4.1.2 for details). The dashed curves are the simulation results for ‘vanilla’ cloud-crushing (in yellow) and cloud-crushing in a CC85 background (in teal). As a specific example which would clearly demonstrate departure from model prediction, we choose the simulation result with $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.8$ and $\mathcal{M}=1.496$ (cf. Table 1; not the fiducial parameter set). The corresponding model predictions are shown with solid lines in the same colour as their dashed counterparts. The thick, solid light-coloured curves use the original FB22 single-cloud model and does not match well with the simulation results. On changing to our ‘modified’ model that floors the inflow velocity according to Eq. 4.1 (described in Section 4.1.2 & Appendix C), we are able to achieve a better match with the simulation results – the solid thin dark coloured curves. For simplicity, the free parameters of the model are chosen by visual inspection and we simply demonstrate the need to construct a robust single-cloud model – a necessary step to explain the evolution of clouds in realistic outflows (cf. Fig. 14). The inset plot shows an additional blue tightly dotted curve up to $100\ t_{\rm cc,ini}$, which corresponds to the same cloud-crushing simulation in a CC85 wind as the other simulation (loosely dashed blue) but run at a higher resolution of $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=16$ instead of $8$. Changes in the cloud mass evolution due to increase in simulation resolution are too minimal to account for the discrepancy between the FB22 prediction and our simulation result.